Show your solution and do not forget the units.

1. Calculate the mole fraction of each components of a solution containing 65 g of ethanol (C2H6O) in 350 g of water.

2.What are the mole fraction big hydrochloric acid (HCl) and water in a 20% (w/w) aqueous HCl solution.


1
Expert's answer
2021-10-15T02:54:18-0400

1) Atomic mass of C, H, O, Cl, Na are 12, 1, 16, 35.5 and 23 respectively.

Mass of ethanol, w1 = 65 g

Molar mass of Ethanol (C2H5OH), M2

= 2×12 + 5×1 + 16 + 1 g/mol

= 46 g/mol

Moles of ethanol, n1 = w1/M1 = 65/46

= 1.413mol

Mass of water, w2 = 350 g

Molar mass of H2O, M2 = 2×1 + 16 = 18g/mol

Moles of water = 350/18 = 19.44 mol

Mole fraction of ethanol = n1/(n1+n2)

= 1.413/(1.413+19.44)

= 1.413/20.86

= 0.93

Sum of mole fraction of components is always 1.

Mole fraction of water = 1- 0.93 = 0.07